15 cubic ft of air at 300 psia and 420°F at constant volume. What are

(a) the final pressure

(b) the work

(c) the change of internal energy

(d) the transferred heat

(e) the change of enthalpy and

(f) the change entropy?


1
Expert's answer
2022-02-10T15:10:09-0500

P=300psia

T=430°F

V=15ft3

(a)

"\frac{P_1}{P_2}=\frac{T_1}{T_2}"

"P_2=221.78psia"

(B) work done by=zero

dV=0

W=P.dV =0J

(3)

Change internal energy

"\u2206U=\u2206Q-PdV"

Put value

"\u2206U=-908.89BTU"

(4)

Transfer heat

"Q=-\u2206U=-908.89BTu"

(5)

∆H=U+P∆V

"\u2206H=-1272.4598BTU"

(6)

Entropy

"∆S=\frac{dQ}{T}"

"ds=-1.2341BTu\/R"